# Beyond Direct Generation: A Decomposed Approach to Well-Crafted Screenwriting with LLMs

### v1

@@ 4.1 Experimental Setting
Datasets. Our experiments utilize two primary datasets: a large-scale dataset for fine-tuning and a custom high-quality dataset for evaluation.
  • Fine-Tuning Dataset: We constructed a comprehensive fine-tuning dataset comprising $50,000$ samples covering diverse genres, such as historical costume drama, fantasy/cultivation, espionage thrillers, and contemporary urban stories. Each sample consists of structured input (outline, previous events, character profiles, etc.) paired with the target output (narrative directives and novel for Stage 1 training, or screenplay for end-to-end training).
  • Evaluation Dataset: Current LLM benchmarks primarily assess general reasoning and instruction-following capabilities, lacking focus on complex creative tasks such as screenplay generation. We then built a high-quality test set tailored for screenplay evaluation. It comprises $32$ distinct scenes sourced from four different television series, carefully curated by human experts to evaluate plot coherence, thematic relevance, and character development.
Evaluation Metrics. Given the creative and subjective nature of screenplay writing, our evaluation relies on expert human judgment, supplemented by comparative and diagnostic metrics. We developed a comprehensive evaluation framework centered on a detailed scoring rubric.
In addition, we designed several auxiliary metrics to enable a more comprehensive analysis of model performance.
Models for Comparison. The experimental design encompasses a diverse set of models for comprehensive evaluation, including both fine-tuned models and proprietary state-of-the-art APIs.
  • Fine-Tuned Models: The core of this investigation involves models fine-tuned on the custom dataset. Several models of varying scales and specializations are compared to analyze their impact on screenplay generation. As baselines, general-purpose chat models of different scales are included: the smaller-scale models Generic-LLM (Qwen-14B-Chat and QwQ-32B), and the larger Qwen-72B-Chat. The primary focus of this comparison is Qwen-72B-CPT, which was created by performing continual pre-training (CPT) on Qwen-72B. The CPT utilized a narrative corpus of approximately 30 billion tokens, curated from high-quality novels and professional screenplays.
  • API Models: For comparison with state-of-the-art API models, the best-performing fine-tuned model is evaluated against Claude-Sonnet-4 and Gemini-2.5-Pro using a carefully crafted prompt that follows the same structured input format.
[…]
@@ 4.2 Main Results
Table 2: Ablation study of the DSR framework. We evaluate the impact of key components including model scale, continual pre-training (CPT), task-decoupled pipeline, and data synthesis strategies. The best score is in bold, and the second-best is underlined.
@@
4.3 Ablation Study and Analysis
To investigate the impact of different base models and data synthesis strategies on final performance, we conduct a comprehensive ablation study, with results presented in Table 2. Our analysis yields the following key findings:
@@ 4.3 Ablation Study and Analysis
These findings collectively highlight the importance of appropriate model scale, domain-adaptive pre-training, and structured data synthesis in advancing LLMs for sophisticated creative applications such as screenplay writing.
5 Conclusions
Limitations
Appendix A Prompts for Data Synthesis
@@
A.1 Prompts for Structured Input (X)
@@
A.2 Prompts for Narrative Directives (Ic)
@@
A.3 Prompts for Enhanced Forward Synthesis (Novel N)
@@
Appendix B Inference Prompts
This section presents the specific prompts used during the inference phase of our DSR framework. These prompts embody the core principle of our approach, explicitly decoupling creative narrative generation from stylistic format conversion. As mentioned in the previous section, placeholders for dynamic inputs are enclosed in curly braces {}, and all experiments were conducted using the original Chinese prompts, not the provided English translations. The prompt for creative narrative generation, shown in Figure 19, takes a scene outline and other structured information as input to generate a rich, novel-style prose output. This narrative prose then becomes the input for the stylistic format conversion prompt, detailed in Figure 21. This prompt’s task is to reformat the text into a professionally formatted screenplay.
@@ Appendix B Inference Prompts
Appendix C Case Study

### v2

@@ 4.1 Experimental Setting
Datasets. Our experiments utilize two primary datasets: a large-scale dataset for fine-tuning and a custom high-quality dataset for evaluation. [^2] See Appendix B.1 for detailed dataset statistics and composition.
  • Fine-Tuning Dataset: We constructed a comprehensive fine-tuning dataset comprising $50,000$ samples covering diverse genres.
  • Evaluation Dataset: We built a high-quality test set tailored for screenplay evaluation, comprising $32$ distinct scenes sourced from four different television series, carefully curated by human experts.
Evaluation Metrics. Given the creative and subjective nature of screenplay writing, our evaluation relies on expert human judgment, supplemented by comparative and diagnostic metrics. [^3] See Appendix B.2 for complete metric definitions and scoring rubric details.
The primary evaluation metric is a holistic quality score assigned by professional Chinese television drama screenwriters. We involved over $20$ experienced screenwriters in the evaluation, with each screenplay independently scored by all of them. The final score for each screenplay is the average of all screenwriters’ ratings. The scoring rubric, illustrated in Figure 3, employs a 12-point scale organized into six quality tiers.
In addition, we designed several auxiliary metrics including Variance, Error Counts, Ratio to Human, and Win Rate to enable a more comprehensive analysis of model performance.
Models for Comparison. The experimental design encompasses a diverse set of models for comprehensive evaluation, including both fine-tuned models and proprietary state-of-the-art APIs. [^4] See Appendix B.3 for complete model specifications.
  • Fine-Tuned Models: Several models of varying scales and specializations are compared, including Generic-LLM (Qwen-14B-Chat and QwQ-32B), Qwen-72B-Chat, and Qwen-72B-CPT.
  • API Models: The best-performing fine-tuned model is evaluated against Claude-Sonnet-4 and Gemini-2.5-Pro.
Implementation Details. To ensure a fair and rigorous comparison, all fine-tuned models were trained on the same dataset using identical hyperparameters. [^5] See Appendix B.4 for complete hyperparameters and hardware specifications.
@@ 4.2 Main Results
4.3 Ablation Study
To validate the design choices of our DSR framework, we conduct comprehensive ablation studies. [^6] See Appendix B.5 for complete ablation results and detailed analysis. The key findings are: (1) Model scale is crucial—larger models significantly outperform smaller variants; (2) The two-stage DSR pipeline yields approximately $2.4$ points improvement over end-to-end generation; (3) Both continual pre-training (CPT) and Chain-of-Thought (CoT) reasoning contribute to performance gains; (4) Our Hybrid data synthesis strategy outperforms the Reverse-only approach, achieving the lowest variance of $0.14$. These findings collectively highlight the importance of appropriate model scale, domain-adaptive pre-training, and structured data synthesis in advancing LLMs for sophisticated creative applications.
5 Conclusions
Limitations
Appendix A DSR Framework Implementation Details
A.1 Data Preprocessing Pipeline
The foundation of our synthesis strategy is a high-quality, pre-processed screenplay corpus. As shown in the lower-left corner of Figure 2, all raw scripts first pass through a standardized cleaning pipeline to ensure data quality and consistency. The pipeline includes four main steps: 1) Script Selection filters for relevant screenplays, 2) Format Standardization unifies diverse writing styles into a single schema, 3) Scene Structuring parses scenes into a consistent data format, and 4) Scene Filtering removes noisy or irrelevant scenes.
Appendix B Experimental Details
B.1 Dataset Details
B.1.1 Fine-Tuning Dataset
We constructed a comprehensive fine-tuning dataset comprising $50,000$ samples covering diverse genres, such as historical costume drama, fantasy/cultivation, espionage thrillers, and contemporary urban stories. Each sample consists of structured input (outline, previous events, character profiles, etc.) paired with the target output (narrative directives and novel for Stage 1 training, or screenplay for end-to-end training).
B.1.2 Evaluation Dataset
Current LLM benchmarks primarily assess general reasoning and instruction-following capabilities, lacking focus on complex creative tasks such as screenplay generation. We then built a high-quality test set tailored for screenplay evaluation. It comprises $32$ distinct scenes sourced from four different television series, carefully curated by human experts to evaluate plot coherence, thematic relevance, and character development.
B.2 Evaluation Metrics Details
B.2.1 Primary Metric: Holistic Quality Score
B.2.2 Auxiliary Metrics
B.3 Model Specifications
B.3.1 Fine-Tuned Models
The core of this investigation involves models fine-tuned on the custom dataset. Several models of varying scales and specializations are compared to analyze their impact on screenplay generation. As baselines, general-purpose chat models of different scales are included: the smaller-scale models Generic-LLM (Qwen-14B-Chat and QwQ-32B), and the larger Qwen-72B-Chat. The primary focus of this comparison is Qwen-72B-CPT, which was created by performing continual pre-training (CPT) on Qwen-72B. The CPT utilized a narrative corpus of approximately 30 billion tokens, curated from high-quality novels and professional screenplays.
B.3.2 API Models
For comparison with state-of-the-art API models, the best-performing fine-tuned model is evaluated against Claude-Sonnet-4 and Gemini-2.5-Pro using a carefully crafted prompt that follows the same structured input format.
B.4 Implementation Details
B.4.1 Training Configuration
[…]
B.4.2 Inference Configuration
B.5 Ablation Study: Complete Results and Analysis
To investigate the impact of different base models and data synthesis strategies on final performance, we conduct a comprehensive ablation study, with results presented in Table 2.
Table 2: Complete ablation study of the DSR framework. We evaluate the impact of key components including model scale, continual pre-training (CPT), task-decoupled pipeline, and data synthesis strategies. The best score is in bold, and the second-best is underlined.
@@ B.5 Ablation Study: Complete Results and Analysis
B.5.1 Impact of Model Scale
@@ B.5.4 Data Synthesis Strategy Comparison
Appendix C Prompts for Data Synthesis
@@
C.1 Prompts for Structured Input (X)
@@
C.2 Prompts for Narrative Directives (Ic)
@@
C.3 Prompts for Enhanced Forward Synthesis (Novel N)
@@
Appendix D Inference Prompts
This section presents the specific prompts used during the inference phase of our DSR framework. These prompts embody the core principle of our approach, explicitly decoupling creative narrative generation from stylistic format conversion. As mentioned in the previous section, placeholders for dynamic inputs are enclosed in curly braces {}, and all experiments were conducted using the original Chinese prompts, not the provided English translations.
The prompt for creative narrative generation, shown in Figure 19, takes a scene outline and other structured information as input to generate a rich, novel-style prose output. This narrative prose then becomes the input for the stylistic format conversion prompt, detailed in Figure 21. This prompt’s task is to reformat the text into a professionally formatted screenplay.
@@ Appendix D Inference Prompts
Appendix E Case Study